OpenStudy (anonymous):

Prove that if |x-x0|

OpenStudy (anonymous):

\[\left| x-x0 \right|<\min(\frac{ ε }{ 2(|y0|+1)},1)\] \[\left| y-y0 \right|<\frac{ ε }{ 2(|x0|+1) }\]

OpenStudy (anonymous):

then \[\left| xy-x0y0 \right|<ε\]

OpenStudy (anonymous):

Hint: \[|xy-x_0y_0|=|xy-x_0y+x_0y-x_0y_0|\]

OpenStudy (anonymous):

|y(x-x0) + x0(y-y0)|<ε it goes like that next right?

OpenStudy (anonymous):

ive reached that...the point is i dont know what to do after that..

OpenStudy (anonymous):

you'll want to apply the triangle equality first:\[|xy-x_0y+x_0y-x_0y_0|\le|xy-x_0y|+|x_0y-x_0y_0|\]Then factor and apply your hypothesis.

OpenStudy (anonymous):

triangle inequality*

OpenStudy (anonymous):

oh oh triangle inequality ahhhh never thought of that thxxx